OpenStudy (anonymous):

Let me post my question right quick. ^_^ Btw, i will give medals. P.s. I just want to know how to solve it, not the answer. ^_^ Thanks a ton!

OpenStudy (anonymous):

\[18r ^{4}s ^{5}t ^{6}\div -3r ^{2}st ^{3}\]

OpenStudy (baru):

\[\frac{a^m}{a^n}=a^{m-n}\]

OpenStudy (baru):

its a direct application of the above rule

OpenStudy (anonymous):

so 18r^4/-3r^2 is r^4-^2 ?

OpenStudy (anonymous):

what would you do with the 18, and -3?

OpenStudy (baru):

\[\frac{18r ^{4}s ^{5}t ^{6}}{-3r ^{2}st ^{3}}\]

OpenStudy (anonymous):

or, is there a way to make it 18rst^?

OpenStudy (anonymous):

and -3rst^?

OpenStudy (baru):

\[\frac{18}{-3} \times \frac{r^4}{r^2} \times \frac{s^5}{s} \times \frac{t^6}{t^3}\]

OpenStudy (anonymous):

oh, i get it.

OpenStudy (baru):

simplify the above using the rule \[\frac{a^m}{a^n}=a^{m-n}\]

OpenStudy (anonymous):

so its -6r^2s^5t^3 ?

OpenStudy (baru):

yep... but the "s" term is wrong

OpenStudy (anonymous):

hw so?

OpenStudy (anonymous):

oh, would it be s^4?

OpenStudy (baru):

yes :)

OpenStudy (anonymous):

oh, so if there isnt a ^? than it is automatically -1 from ^5

OpenStudy (anonymous):

?

OpenStudy (baru):

any number raised to the power "1" is the number itself... for example \(s=s^1\)

OpenStudy (anonymous):

just as i suspected. ok, thanks! ^_^

OpenStudy (baru):

sure :)